OpenStudy (anonymous):

A box contains four 40-W bulbs, five 60-W bulbs, and six 75-W bulbs. If bulbs are se- lected one by one in random order, what is the probability that at least two bulbs must be selected to obtain one that is rated 75W?

OpenStudy (anonymous):

so first one is not a 75W bulb: P(not a 75Wbulb)=(4+5)/(4+5+6)=9/15 P(a 75W bulb)=6/(15-1)=6/14 9/15*6/14=9/35